A yardstick measures 0.5 by 2 by 36 inches. How many yardsticks will fit in a box 3.5 inches wide and 36 inches high, if the length of the box is 18 inches?
Answer by checkley77(12844)

.5*2=1 in^2 is the end area of the yard sticks.
3.5*18=63 in^2 is the area of the box end
63/1=63 yard sticks can fit in this box.
Proof:
3.5/.5=7
18/2=9
7*9=63 sticks will fit.